Queue Observing Issues

Monday, January 30, 2017 1:23 PM

Problem(s) Encountered:

Dome would randomly jump ~180 during exposures, went out noted dome position and reset encoders .  This worked however it needed to be repeated after about two exposures.  

Really confused with what happened the rest of the night...

Had already taken spectra with no problems.  The guider computer was set up so screen had a box roughly where the guide star needed to be in order for target to be in slit.    

Guide star did not show up and quick search nearby showed nothing. Tried to move around to different guide stars but this did not work.  Reset MIS and tried again.  Thought this fixed the problem but on exposure could not locate target.  Tried cycling MIS one more time

Next guess was the telescope was in the wrong spot, double checked RA& Dec listed.  Went to nearest brightest star and it was not there (had done this at start of the night with no problems).  Went to zenith and did TCS reboot, realigned with tilt sensors.  Observed another target, everything worked fine

Next target up in rotator angle, target guider was not going to correct location again,  thought this was an issue with the rotator angle (have had issue before) reset guider/MIS did not help.  Moved rotator back and tried another target and believe it was the same problem as above


Double checked all RA & Dec were listed correctly and that inner slit was selected in Jskycal and tracking rates/ rotator angle were correct


Not sure what happened tonight,  Went back and looked at images where I could not find target, found one , ran

osctrtask, it asked for + 2500dx

Not sure what happened or why it failed so quickly after TCS reset


Not really sure what is being reported here.  Was the problem that the target was seen in the science camera but no guide star was found, or was everything off, indicating a telescope set issue?

I can say that the tilt sensors appeared to be off some relative to what they should read at zenith (since replacing the load cell on HP2, tilt sensors have been around 0.11 & 0.07 when at zenith).

I went ahead and brought everything down and back up (excepting the HE and IE on OSMOS, which remained powered up).  It was worthwhile to give hiltner and the workstation a reboot anyways.  With everything fresh, I’ve been moving the telescope around, checking the guider (as much as one can during the day) and checking the MIS.  All appears well to me.  We’ll see what happens tonight.  Just please remember to set to the appropriate tilt sensor values if reseting encoders at zenith.

As for the dome, I went ahead and replaced the chip set.  I did this after seeing a similar sort of jump in encoder readout.  I also “wiggled” the wires to the IR heads and blew them off with compressed air.  After doing this, the system appears to be working a bit better.  Not sure yet how much better.  Only time will tell.  

One potential way to keep the dome from walking away in the middle of an observation is to turn “Auto Dome” off during scans.  As error builds up, auto dome could be briefly turned back on to zero the error (error > 5 degrees or so) after ensuring that the readout for dome azimuth is roughly equal to that of the telescope.  Once the dome resets to match telescope azimuth, auto dome could then be turned back off.  In the meantime, if the dome azimuth readout changes to a nonsensical value, the observer could reset it through the TCS GUI before enabling auto dome.  This would certainly require more user interface and tending, but it might be useful if desired.